• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 22:38:03 on July30,2003

    Entry number 108224

    keyword=Turn on septum


    The septum has reached below 7 K on both upper and lower coils.

    He Level is 87%

    He Gas Flow L1 and L2 are both 11.9 SLM.

    Interlocks reset happily.

    Dump Switch CMD ON (light was already green)

    Power Supply CMD ON (light turns green)

    Choose set current of 135 A (we will need 270.6A at this next
    kinematic, but I am told to do this stepwise).

    Current starts to ramp up.

    Below is the stripchart for the cooldown.




    FIGURE 1